OpenStudy (anonymous):

\[ \frac{dy}{dx}=\sqrt{y^2-1} \]

OpenStudy (anonymous):

Boundary conditions are\[y(a)=y(-a)=1\]

myininaya (myininaya):

\[\frac{dy}{\sqrt{y^2-1}}=dx\] Separation of variables. Now integrate both sides. :

OpenStudy (anonymous):

Why didn't I see that?! Thanks